### 13 Dec 2023 v2

# Enabling Fast 2-bit LLM on GPUs: Memory Alignment and Asynchronous Dequantization

Abstract
Large language models (LLMs) have demonstrated impressive abilities in various domains while the inference cost is expensive. Many previous methods use 2-bit quantization to reduce LLM inference cost by reducing storage and accelerating computation. However, challenges still exist in reducing LLM inference cost: (1) Large accuracy loss for 2-bit quantization. Weights are quantized by groups, while the gradient of weight are large in some groups, resulting in large quantization errors and accuracy loss. (2) Time-consuming dequantization operations on GPUs. Mainstream methods require a dequantization operation to perform computation on the quantized weights, and the 2-order dequantization operation is applied because scales of groups are also quantized. These dequantization operations lead to >50% execution time, hindering the potential of reducing LLM inference cost.
To tackle these challenges, we propose the following techniques in this paper. (1) Range-aware quantization with memory alignment. We only quantize a small fraction of groups with the larger range using 4-bit with memory alignment consideration on GPUs. (2) Asynchronous dequantization. We point out that calculating the scales of each group is independent of the loading weights of each group. Thus, we design the asynchronous dequantization on GPUs, leading to up to 3.92 speedup. Our method achieves 2.85-bit for each weight considering all scales/zeros for different models. The end-to-end speedup for Llama2-7b is 1.74 over the original model.
1 Introduction
Large language models (LLMs) demonstrate remarkable capabilities in various domains like natural language understanding and generation [1, 2]. However, their computational expense during inference is noteworthy. For example, OpenAI reveals that the GPT-4 Turbo incurs a cost of millions of dollars each day [3][4]. Many previous studies exploit quantization methods to reduce LLM inference cost by reducing storage and accelerating computation [5, 6, 7, 8, 9]. The state-of-the-art methods have realized 2-bit quantization for mainstream LLMs, exemplified by the Llama-2 family [10]. These methods quantize weights by group to improve accuracy and the scales are also quantized (2-order) to further reduce average bit [11].
However, there are still existing challenges of 2-bit quantization. (1) Large accuracy loss for 2-bit quantization. Both the post-training quantization (PTQ, e.g., GPTQ [5]) and the quantization-aware training (QAT, e.g., LLM-QAT [12] and Greenbit [11]) approaches incur the 3.2% to 5.6% accuracy loss for Llama2-7b with 2-bit quantization [5, 11]. (2) Time-consuming dequantization operations on GPUs. Mainstream methods require a dequantization operation to perform computation on 2-bit weights. Furthermore, a 2-order dequantization operation is applied because the scales of groups are also quantized. The dequantization operation takes over 50% of the total execution time. These challenges become the crucial factors impeding fast and low-cost LLM inference.
In response to these challenges, we enable fast and low-cost LLM inference on GPUs with the following contributions in this paper:
Range-aware quantization with memory alignment. We point out that the range of weights by groups varies. Thus, only 25% of the weights are quantized using 4-bit with memory alignment. Such a method reduces the accuracy loss for 2-bit Llama2-7b quantization from 8.7% to 2.9%.
Asynchronous dequantization. We point out that calculating the scales of each group is independent of the loading group weights. Thus, we design the asynchronous dequantization and accelerate the GPU kernel by up to 3.92.
We conduct extensive experiments on Llama-2 model families. We achieve 2.85-bit for each weight considering all scales/zeros for different models. The end-to-end speedup for Llama2-7b is 1.74, and we reduce both runtime cost and hardware cost by up to 2.70 and 2.81 with less GPU requirements.
2 Method
2.1 Range-aware Quantization
Challenge. Quantization uses discrete values to represent continuous weights, leading to quantization errors. Applying 2-bit quantization leads to larger quantization errors and further results in larger accuracy loss for Llama2-7b. Even the state-of-the-art 2-bit methods still suffer from >3% accuracy loss for Llama2-7b [5, 11].
Motivations and Insights. Some previous methods like SpQR [7] and AWQ [6] have analyzed that the quantization errors are large in specific groups. Applying 4-bit quantization to these groups with a large range can reduce the quantization error. Thus, our key insight is, only a small fraction of weights requires 4-bit quantization and average bit can be still reduced by applying 2-bit quantization on the other groups.
Approach: 2/4-bit Quantization. In Figure 1, we depict the quantization flow of weights. First, we analyze the range distribution by group. The range of groups in the same input channels is merged to streamline the complexity of hardware computation. Thus, the range of weights by groups is equivalent to the amplitude variance in the grouped input channels. Then, we analyze the variance and obtain the mixed-precision layout that the input channels with large variance require 4-bit quantization and others require 2-bit quantization. According to the mixed-precision layout, we apply the finetuning approach and then quantize different grouped input channels with different bit-width. For small range input channels, we quantize the weights with 2-bit and also quantize the scales with 4-bit to reduce average bit. For large range input channels, we quantize the weights with 4-bit. Equation 1 is the average bit of 2/4-bit quantization.
where is proportion of weights quantized with 2-bit, and are the group sizes of 1-order and 2-order quantization separately, is the bit precision of 2-order quantization.
Approach: Memory Alignment. As illustrated in Figure 1, the quantized weights are split into two quantized weight matrixes. The first matrix stores 3 groups of 2-bit quantized weights (the size of group is 16) and 8 4-bit quantized weights into continuous memory. The other matrix stores 8 remaining 4-bit quantized weights into independent memory. Thus, the memory access for each thread in GPU is memory-aligned. Then, the zeros and scales are packed together and two of three scales are compressed to fill 16-bit memory. These memory alignment approaches preserve accuracy loss and only lead to <0.02 average bit overheads.
2.2 Asynchronous Dequantization
Challenge. Before the multiplication operations, the quantized weights are required to be dequantized to 16-bit. Previous designs (e.g., SPQR[7]) use the synchronous dataflow (i.e., performing dequantization after loading all quantized weights, scales and zeros), resulting in >50% overheads of end-to-end execution time. Thus, the key challenge is the time-consuming synchronous dequantization operation becomes the bottleneck in accelerating LLM inference after quantization.
Motivation and Insights. The 1-order and 2-order quantization requires two synchronous dequantizations in dataflow. However, the 2-order dequantization for calculating the scales of each group is independent of the weights of each group for the 1-order dequantization. Moreover, the 1-order dequantization for restoring the weights to 16-bit is independent of the input data of multiplication. Thus, our key insight is calculating the scales of each group can be overlapped with loading weights of each group in GPU kernel.
Approach: Asynchronous dequantization. Based on the insights above, we design the asynchronous dequantization dataflow as illustrated in Figure 2 on GPUs. With the help of the shared memory, we can overlap the calculating scales of each group and the loading weights of each group. Further, we use CUDA primitive __shfl_down_sync() to reduce the result inside a warp efficiently. From the perspective of memory access, we apply vector load (e.g. double4) technique to load quantized weights, activations, scales and zeros to minimize the numbers of memory access.
3 Experiments
We conduct comprehensive experiments on the Llama2-7b model family [10]. The accuracy of the introduced quantized LLMs is assessed across four zero-shot benchmarks, namely PIQA [13], HellaSwag [14], WinoGrande [15], ARC-Easy [16]. We use 2-order weight-only quantization to evaluate the performance.
Baselines. We compare our design with state-of-the-art quantization designs, including GPTQ [5] and Greenbit [11]. The original PyTorch implementation on HuggingFace [17] is used as the baseline.
Hardware Platforms. We implement our design and compare other baselines on NVIDIA RTX 2080, NVIDIA RTX 3090, and NVIDIA A100 (80G) GPUs with CUDA version 12.2. The runtime cost is evaluated considering the power consumption using the NVML library [18], and the hardware cost is evaluated considering the price ($12,500 for A100, $1499 for RTX 3090, and $699 for RTX 2080) with the 5-year service life period.
| Method | PIQA | HellaSwag | WinoGrande | ARC-Easy | Average | |
| Llama2-7b | 16 | 77.2% | 56.1% | 68.7% | 70.2% | 68.0% |
| GPTQ | 3 | 71.7% | 48.2% | 61.2% | 56.1% | 59.3% |
| Greenbit | 2.91 | 77.2% | 53.8% | 65.8% | 62.5% | 64.8% |
| Ours | 2.85 | 75.9% | 51.3% | 66.4% | 66.9% | 65.1% |
| Llama2-13b | 16 | 78.8% | 59.7% | 69.6% | 73.3% | 70.4% |
| GPTQ | 3 | 75.2% | 56.1% | 64.3% | 64.2% | 64.0% |
| Ours | 2.85 | 76.9% | 55.6% | 66.2% | 69.2% | 67.0% |
| Model | Method | GPU | Power(W) | Token/s | End-to-end speedup | Runtime cost reduction | Hardware cost reduction |
| Llama2-7b | PyTorch | 30901 | 290/350 | 25.9 | 1 | 1 | 1 |
| GPTQ | 30901 | 320/350 | 37.5 | 1.45 | 1.31 | 1.39 | |
| Our | 30901 | 350/350 | 45.2 | 1.74 | 1.44 | 1.60 | |
| Our | 20801 | 150/215 | 34.0 | 1.31x | 2.53 | 2.81 | |
| PyTorch | 30902 | 240/350 | 22.0 | 1 | 1 | 1 | |
| Llama2-13b | GPTQ | 30901 | 270/350 | 23.1 | 1.06 | 1.88 | 2.02 |
| Our | 30901 | 250/350 | 25.0 | 1.14 | 2.19 | 2.24 | |
| PyTorch | A1002 | 400/400 | 38.5 | 1 | 1 | 1 | |
| Llama2-70b | GPTQ | A1001 | 400/400 | 46.5 | 1.21 | 2.42 | 2.42 |
| Our | A1001 | 400/400 | 50.5 | 1.31 | 2.62 | 2.62 |
Accuracy Evaluations. Table 1 shows algorithm accuracy with zero-shot performance on Llama2 family. We set the group size to 16 for Llama2-7b and Llama2-13b. Compared with the state-of-the-art GPTQ [5] method, our design significantly reduces the accuracy loss from 8.7% to 2.9% for Llama2-7b, and from 6.4% to 3.4% Llama2-13b. Compared with Greenbit [11] on Llama2-7b, we achieve 0.3% higher accuracy with the lower average bit.
Kernel Speedup. We compare our kernel performance with some mainstream quantized kernels(e.g., GPTQ [5]) and the original FP16 models in common LLM GEMV cases. Our kernel achieves over FP16 on various NVIDIA GPUs.
Performance Evaluations. We compare the inference cost (end-to-end speedup, runtime cost, and hardware cost) among our method, PyTorch, and GPTQ in various model families and model sizes in Table 2. All results are normalized to PyTorch. In the Llama2 model family, our design achieves end-to-end speedup and far outperforms other previous works (e.g., GPTQ[5]) in terms of running cost reduction, reaching over the original model. Moreover, the hardware requirement is dramatically lowered in our design for all models in Table 2, bringing up to runtime cost reduction and hardware cost reduction. The performance underscores the effectiveness of our design’s optimizations and the practical utility of our approach tailored to optimize the inference and deployment of LLMs.